The kinetic coefficient friction of a rubber tire on wet concrete is 0.70. What is the mass of a tractor that with its brakes locked, can be dragged by a tow truck exerting a horizontal force of 1.0 x104 N?
1
Expert's answer
2019-11-01T11:41:07-0400

The friction force with increasing horizontal 'drag' force will be increase too until it reaches the maximum value (Amonton-Coulomb law)


"\left| {{{\vec F}_{{\text{fr}}{\text{.max}}}}} \right| = \mu N"


where "N" - the absolute value of normal force, by the third Newton's law equal to the gravitational force "N = mg"

Then the condition that the tractor can be dragged by a truck is


"F > \left| {{{\vec F}_{{\text{fr}}{\text{.max}}}}} \right| = \mu mg"

Thus


"m < \frac{F}{{\mu g}}"

Using "g \approx 9.8[\frac{{\text{m}}}{{{{\text{s}}^{\text{2}}}}}]" we've got


"m < \frac{{1 \cdot {{10}^4}[{\text{N}}]}}{{0.7 \cdot 9.8[\frac{{\text{m}}}{{{{\text{s}}^{\text{2}}}}}]}}"

And


"m < 1457.7[{\text{kg}}]"